# A Comprehensive Characterization of Galaxy-Cool CGM Connections at $z<0.4$ with DESI Year 1 Data

### v1

@@ 5 Conclusions
    For star-forming galaxies, the absorption strengths correlate with the stellar mass of galaxies from $10^{8}\,M_{\odot}$ to $10^{11}\,M_{\odot}$ with $\propto M_{*}^{0.5}$, while such a trend is weaker for quiescent galaxies. With a fixed stellar mass, we also find a positive correlation between absorption strengths and SFR for star-forming galaxies, following $\propto\mathrm{SFR}^{0.3}$. In addition, for $>10^{10}\,M_{\odot}$ galaxies, the absorption strengths are stronger around star-forming galaxies than those around quiescent galaxies within $30\,\mathrm{kpc}$. The correlations between Ca II and galaxy properties are detected at $z\sim 0.1$ and $z\sim 0.3$.
@@
    After normalizing the impact parameter by the virial radius, the absorption profiles around galaxies with different stellar masses from $10^{8}\,M_{\odot}$ to $10^{11}\,M_{\odot}$ become consistent with each other, indicating that the cool CGM distributes similarly in the halos and is closely linked to halo masses. For quiescent galaxies with masses $>10^{11}\,M_{\odot}$, the Ca II absorption in the inner CGM appears to be weaker. This halo mass trend is also observed at higher redshifts, suggesting that the underlying mechanism(s) regulating such a property persists across cosmic time.
@@
Our results provide a comprehensive characterization of galaxy-cool CGM connections below $z<0.4$ down to $10^{8}\,M_{\odot}$ galaxies, offering novel constraints on the models of galaxy evolution. Together with similar measurements at higher redshifts, the relationships between the properties of galaxies and the cool CGM have been mapped across $\sim 10$ Gyr from redshift 0 to redshift $\sim 2$. Understanding the physical mechanisms behind these relationships over cosmic time will be essential for advancing the knowledge of galaxy formation and evolution.
The analysis adopted in this work can be directly applied to upcoming datasets. For example, at the local Universe, the complete 5-year DESI dataset is expected to further enhance the S/N of the similar measurements by at least a factor of two. In addition, one can combine the images provided by Euclid (Euclid Collaboration et al., 2024) to obtain more precise measurements of gas distribution as a function of azimuthal angle, a key constraint on how gas flows around galaxies. Similar measurements will also be obtained at even higher redshifts. The Prime Focus Spectrograph (PFS) survey (Takada et al., 2014) and the DESI-II (Schlegel et al., 2022), for example, will observe millions of galaxies at $z>2$, enabling explorations of galaxy-CGM connections with large datasets at the peak of the star-formation density of the Universe (e.g., Madau & Dickinson, 2014). Similar to the properties of galaxies, such as stellar mass functions and luminosity functions at different redshifts (e.g., Kelvin et al., 2014; Driver et al., 2022; Weaver et al., 2023), precise measurements of gas around galaxies at different redshifts will serve as fundamental constraints, leading to a more comprehensive understanding of the cosmic baryon cycle.
@@
All data points shown in the figures are available at Zenodo doi: 10.5281/zenodo.14970325.
@@
We thank Siwei Zou and Abhijeet Anand for their constructive comments that greatly helped improve the paper during the DESI Collaboration Wide Review. YVN, TWL, and YLC acknowledge supports from the National Science and Technology Council (MOST 111-2112-M-002-015-MY3, NSTC 113-2112-M-002-028-MY3), the Yushan Fellow Program by the Ministry of Education (MOE) (NTU-110VV007, NTU- [^1] V1007-2, NTU-112V1007-3, NTU-113V1007-4), National Taiwan University research grant (NTU-CC- [^1] L894806, NTU-CC-112L893606, NTU-CC-113L891806, NTU- [^1] L7318, NTU-112L7302). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall z-band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys was supported by: the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
@@ Appendix A Cloudy simulation
We perform Cloudy simulations (Chatzikos et al., 2023) to evaluate the dust depletion factor in the CGM. The simulations are initialized with the Haardt & Madau (2012) extragalactic radiation background at $z=0.2$, and run over a hydrogen volume density range of $10^{-2}$ to $10^{0.5}\,\mathrm{cm^{-3}}$ until the neutral hydrogen column density reaches $10^{19}$, $10^{19.3}$ and $10^{19.5}\,\mathrm{cm^{-2}}$. In order to determine the depletion factor, we run two sets of simulations. First, we run a simulation by adopting the default solar composition in Cloudy (Grevesse & Sauval, 1998; Allende Prieto et al., 2001; Holweger, 2001; Allende Prieto et al., 2002) to obtain the intrinsic column density ratio of Ca II and Zn II (the left panel of Figure 12). Next, we reduce the calcium abundance to simulate dust depletion, adjusting it to match the observed average Ca II-to-Zn II column density ratio of $\sim 0.86$ from Lan & Fukugita (2017) (the middle panel of Figure 12). By comparing the intrinsic ratio with the depleted simulations, we derive an average Ca depletion factor of $\sim 89$ and adopt this value to correct the depletion effect.
Figure 12: Cloudy simulation results across a total hydrogen volume density range of $10^{-2}$ to $10^{0.5}\,\mathrm{cm^{-3}}$ for $\log_{10}{(N_{\mathrm{HI}}/\mathrm{cm^{-2}})}=19.0,19.3,19.5$. The left panel shows the intrinsic Ca II-to-Zn II column density ratio, assuming an initial solar abundance. The middle panel shows the observed ratio from Lan & Fukugita (2017) (dashed line) and simulated ratios by reducing the calcium abundance by a factor of $10^{-1.95}$. The right panel presents the depletion factor, obtained by dividing the intrinsic ratio from the left panel by the simulated depleted values from the middle panel.
@@ Appendix B Comparison with SDSS
Here we compare our DESI measurements with the SDSS measurements from Zhu & Ménard (2013b). To match the SDSS galaxy population, we select DESI galaxies with $r<17.7$ (Strauss et al., 2002). As shown in Figure 13, the DESI measurements are consistent with the SDSS measurements. We note that while DESI provides more galaxy-quasar pairs, the typical S/N of individual DESI quasar spectra is lower than the S/N of SDSS quasar spectra. Together, these two factors yield similar S/N values of the SDSS and DESI composite spectra. Nevertheless, the DESI dataset enables us to probe the Ca II absorption around galaxies with lower mass and higher redshifts.
Figure 13: Comparison of Ca II absorption (K line) around $r<17.7$ galaxies with SDSS and DESI data. The grey data points show the measurements from Zhu & Ménard (2013b) and the blue data points show the measurements obtained in this work with DESI data.

### v2

@@ 5 Conclusions
    For star-forming galaxies, the absorption strengths correlate with the stellar mass of galaxies from $10^{8}\,M_{\odot}$ to $10^{11}\,M_{\odot}$ with $\propto M_{*}^{0.5}$, while such a trend is weaker for quiescent galaxies. With a fixed stellar mass, we also find a positive correlation between absorption strengths and SFR for star-forming galaxies, following $\propto\mathrm{SFR}^{0.3}$. In addition, for $10^{10}\,M_{\odot}$ galaxies, the absorption strengths are stronger around star-forming galaxies than those around quiescent galaxies within $30\,\mathrm{kpc}$. The correlations between Ca II and galaxy properties are detected at $z\sim 0.1$ and $z\sim 0.3$.
@@
    After normalizing the impact parameter by the virial radius, the absorption profiles around galaxies with different stellar masses from $10^{8}\,M_{\odot}$ to $10^{11}\,M_{\odot}$ become consistent with each other, indicating that the cool CGM distributes similarly in the halos and is closely linked to halo masses. For quiescent galaxies with masses $10^{11}\,M_{\odot}$, the Ca II absorption in the inner CGM appears to be weaker. This halo mass trend is also observed at higher redshifts, suggesting that the underlying mechanism(s) regulating such a property persists across cosmic time.
@@
Our results provide a comprehensive characterization of galaxy–cool CGM connections below $z<0.4$ down to $10^{8}\,M_{\odot}$ galaxies, offering novel constraints on the models of galaxy evolution. Together with similar measurements at higher redshifts, the relationships between the properties of galaxies and the cool CGM have been mapped across $\sim 10$ Gyr from redshift 0 to redshift $\sim 2$. Understanding the physical mechanisms behind these relationships over cosmic time will be essential for advancing the knowledge of galaxy formation and evolution.
The analysis adopted in this work can be directly applied to upcoming data sets. For example, at the local Universe, the complete 5 yr DESI data set is expected to further enhance the S/N of the similar measurements by at least a factor of two. In addition, one can combine the images provided by Euclid (Euclid Collaboration et al., 2024) to obtain more precise measurements of gas distribution as a function of azimuthal angle, a key constraint on how gas flows around galaxies. Similar measurements will also be obtained at even higher redshifts. The Prime Focus Spectrograph survey (Takada et al., 2014) and the DESI-II (Schlegel et al., 2022), for example, will observe millions of galaxies at $z>2$, enabling explorations of galaxy–CGM connections with large data sets at the peak of the star formation density of the Universe (e.g., Madau & Dickinson, 2014). Similar to the properties of galaxies, such as stellar mass functions and luminosity functions at different redshifts (e.g., Kelvin et al., 2014; Driver et al., 2022; Weaver et al., 2023), precise measurements of gas around galaxies at different redshifts will serve as fundamental constraints, leading to a more comprehensive understanding of the cosmic baryon cycle.
@@
All data points shown in the figures are available at Zenodo DOI:https://doi.org/10.5281/zenodo.17083730 (catalog 10.5281/zenodo.17083730).
@@
We thank the anonymous referee for the constructive report. We also want to thank Siwei Zou and Abhijeet Anand for their constructive comments that greatly helped improve the paper during the DESI Collaboration Wide Review. Y.V.N., T.W.L., and Y.L.C. acknowledge support from the National Science and Technology Council (MOST 111-2112-M-002-015-MY3, NSTC 113-2112-M-002-028-MY3), the Yushan Fellow Program by the Ministry of Education (MOE) (NTU-110VV007, NTU- [^1] V1007-2, NTU-112V1007-3, NTU-113V1007-4) and National Taiwan University research grant (NTU-CC- [^1] L894806, NTU-CC-112L893606, NTU-CC-113L891806, NTU- [^1] L7318, NTU-112L7302). This material is based upon work supported by the U.S. Department of Energy (DOE), Office of Science, Office of High-Energy Physics, under Contract No. DE–AC02–05CH11231, and by the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility under the same contract. Additional support for DESI was provided by the U.S. National Science Foundation (NSF), Division of Astronomical Sciences under Contract No. AST-0950945 to the NSF’s National Optical-Infrared Astronomy Research Laboratory; the Science and Technology Facilities Council of the United Kingdom; the Gordon and Betty Moore Foundation; the Heising-Simons Foundation; the French Alternative Energies and Atomic Energy Commission (CEA); the National Council of Humanities, Science and Technology of Mexico (CONAHCYT); the Ministry of Science and Innovation of Spain (MICINN), and by the DESI Member Institutions: https://www.desi.lbl.gov/collaborating-institutions. The DESI Legacy Imaging Surveys consist of three individual and complementary projects: the Dark Energy Camera Legacy Survey (DECaLS), the Beijing-Arizona Sky Survey (BASS), and the Mayall $z$-band Legacy Survey (MzLS). DECaLS, BASS and MzLS together include data obtained, respectively, at the Blanco telescope, Cerro Tololo Inter-American Observatory, NSF’s NOIRLab; the Bok telescope, Steward Observatory, University of Arizona; and the Mayall telescope, Kitt Peak National Observatory, NOIRLab. NOIRLab is operated by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the National Science Foundation. Pipeline processing and analyses of the data were supported by NOIRLab and the Lawrence Berkeley National Laboratory. Legacy Surveys also uses data products from the Near-Earth Object Wide-field Infrared Survey Explorer (NEOWISE), a project of the Jet Propulsion Laboratory/California Institute of Technology, funded by the National Aeronautics and Space Administration. Legacy Surveys were supported by the Director, Office of Science, Office of High Energy Physics of the U.S. Department of Energy; the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility; the U.S. National Science Foundation, Division of Astronomical Sciences; the National Astronomical Observatories of China, the Chinese Academy of Sciences and the Chinese National Natural Science Foundation. LBNL is managed by the Regents of the University of California under contract to the U.S. Department of Energy. The complete acknowledgments can be found at https://www.legacysurvey.org/. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the U. S. National Science Foundation, the U. S. Department of Energy, or any of the listed funding agencies. The authors are honored to be permitted to conduct scientific research on Iolkam Du’ag (Kitt Peak), a mountain with particular significance to the Tohono O’odham Nation.
@@ Appendix A Cloudy simulation
We perform Cloudy simulations (Chatzikos et al., 2023) to evaluate the dust depletion factor in the CGM. The simulations are initialized with the Haardt & Madau (2012) extragalactic radiation background at $z=0.2$, and run over a hydrogen volume density range of $10^{-2}$$10^{0.5}\,\mathrm{cm^{-3}}$ until the neutral hydrogen column density reaches $10^{19}$, $10^{19.3}$, and $10^{19.5}\,\mathrm{cm^{-2}}$. In order to determine the depletion factor, we run two sets of simulations. First, we run a simulation by adopting the default solar composition in Cloudy (Grevesse & Sauval, 1998; Allende Prieto et al., 2001, 2002; Holweger, 2001) to obtain the intrinsic column density ratio of Ca II and Zn II (the left panel of Figure 12). Next, we reduce the calcium abundance to simulate dust depletion, adjusting it to match the observed average Ca II-to-Zn II column density ratio of $\sim 1.07$ from (Lan & Fukugita 2017; the middle panel of Figure 12). The value is averaged from the Ca II doublet values. By comparing the intrinsic ratio with the depleted simulations, we derive an average Ca depletion factor of $\sim 74$ and adopt this value to correct the depletion effect.
Figure 12: Cloudy simulation results across a total hydrogen volume density range of $10^{-2}$$10^{0.5}\,\mathrm{cm^{-3}}$ for $\log_{10}{(N_{\mathrm{HI}}/\mathrm{cm^{-2}})}=19.0,19.3,19.5$. The left panel shows the intrinsic Ca II-to-Zn II column density ratio, assuming an initial solar abundance. The middle panel shows the observed ratio from (Lan & Fukugita 2017; dashed line) and simulated ratios by reducing the calcium abundance by a factor of $10^{-1.87}$. The right panel presents the depletion factor, obtained by dividing the intrinsic ratio from the left panel by the simulated depleted values from the middle panel.
@@ Appendix B Comparison with SDSS
Here we compare our DESI measurements with the SDSS measurements from Zhu & Ménard (2013b). To match the SDSS galaxy population, we select DESI galaxies with $r<17.7$ (Strauss et al., 2002). As shown in Figure 13, the DESI measurements are consistent with the SDSS measurements. We note that while DESI provides more galaxy–quasar pairs, the typical S/N of individual DESI quasar spectra is lower than the S/N of SDSS quasar spectra. Together, these two factors yield similar S/N values of the SDSS and DESI composite spectra. Nevertheless, the DESI data set enables us to probe the Ca II absorption around galaxies with lower mass and higher redshifts.
Figure 13: Comparison of Ca II absorption (K line) around $r<17.7$ galaxies with SDSS and DESI data. The gray data points show the measurements from Zhu & Ménard (2013b) and the blue data points show the measurements obtained in this work with DESI data.